OpenStudy (joeking):

Tom and his best friend Edward are seeking to join a gym. Tom saw on television that Platinum Gym was having a grand opening in a few months. Tom is considering a membership at Platinum Gym. Platinum Gym charges a one-time registration fee of $60, and membership is $20 a month. Tom's best friend Edward says he plans to join Super Fit because Super Fit only charges $10 a month but has a one-time registration fee of $150. (a)Write an expression to represent the membership cost of Platinum Gym for x months. (b)Write an expression to represent the membership cost of Super Fit for x months.

OpenStudy (michaelbp):

60 + 20(1) = 60 + 20 = 80 per month for Tom at Platinum gym 150 + 10(1) = 150 + 10 = 160 per month for Edward at Super fit so for 1 month, Tom pays less for a year.. Tom pays : 80 * 12 = 960 Edward pays : 160 * 12 = 1920 Tom still pays less...even at 1 yr ======================== 60 + 20m = 150 + 10m 20m - 10m = 150 - 60 10m = 90 m = 9 --- they would pay the same at 9 months 60 + 20(9) = 60 + 180 = 240 is what they would both pay ========================= The better deal depends on how long they want the membership. If the membership is LESS then 9 months, then Tom has the better deal at Platinum..but if the membership is MORE then 9 months, then Edward has the better deal at Superfit.